### v1

# The Mind in the Machine: A Survey of Incorporating Psychological Theories in LLMs

Abstract
Psychological insights have long shaped pivotal NLP breakthroughs, including the cognitive underpinnings of attention mechanisms, formative reinforcement learning, and Theory of Mind-inspired social modeling. As Large Language Models (LLMs) continue to grow in scale and complexity, there is a rising consensus that psychology is essential for capturing human-like cognition, behavior, and interaction. This paper reviews how psychological theories can inform and enhance stages of LLM development, including data, pre-training, post-training, and evaluation&application. Our survey integrates insights from cognitive, developmental, behavioral, social, personality psychology, and psycholinguistics. Our analysis highlights current trends and gaps in how psychological theories are applied. By examining both cross-domain connections and points of tension, we aim to bridge disciplinary divides and promote more thoughtful integration of psychology into future NLP research.
1 Introduction
As Large Language Models (LLMs) grow in scale and complexity, the Natural Language Processing (NLP) community is increasingly recognizing psychology as essential for capturing human-like cognition, behavior, and interaction (Qu et al., 2024; Lewis, 2025). Psychology, grounded in empirically validated and computationally adaptable frameworks (Sartori & Orrù, 2023; Ong, 2024), can address core LLM challenges such as reasoning fidelity, context retention, and user interaction. Reflecting these strengths, psychological insights have long shaped pivotal NLP advances, including the cognitive inspirations of attention mechanisms, formative reinforcement learning approaches, and Theory of Mind-inspired social modeling.
Despite extensive multidisciplinary efforts, a holistic review systematically integrating psychology across the LLM lifecycle remains missing. Most surveys and position papers remain fragmented, typically falling into three broad categories: (1) Some investigate how LLMs can empower traditional psychology or cognitive science research, for instance by modeling human reasoning and behavior at scale (Demszky et al., 2023; Abdurahman et al., 2024; Ong, 2024; Ke et al., 2024). (2) Others approach LLMs as subjects of psychological analysis, aiming to adapt or extend psychological theory – such as personality or cognition frameworks – to interpret and evaluate model behavior (Li et al., 2024b; Hagendorff et al., 2023; tse Huang et al., 2024; Pellert et al., 2024). (3) Finally, a third group leverages a single or limited set of psychological constructs to enhance model alignment or multi-agent frameworks – improving system reliability, social interaction, and trustworthiness (Liu et al., 2023; Dong et al., 2024b). This includes research on social influence for AI safety (Zeng et al., 2024a), moral reasoning in legal tasks (Almeida et al., 2024), and partial integrations of social or developmental psychology (Sartori & Orrù, 2023; Zhang et al., 2024c; Serapio-García et al., 2025). However, no existing work provides a unified map of how diverse psychological sub-areas can be harnessed, from data through application. Our survey fills this gap by offering a stage-wise view of how psychology can strengthen LLM capabilities and alignment across the entire lifecycle.
To address this gap, we present a structured review that situates psychological theories from six major areas across the entire LLM development pipeline. Drawing on research from multiple NLP and AI conferences [^1] We survey 175 papers from major *CL venues (ACL Anthology), plus COLING, NeurIPS, ICML, ICLR, and influential arXiv preprints from late 2021 to early 2025 via keyword search., our contributions are twofold: (1) We systematically review psychological theories applied in key stages of LLM development, identifying gaps and inconsistencies. (2) We highlight under-explored concepts alongside critical issues and debates at the intersection of psychology and NLP. Collectively, these contributions demonstrate how integrating diverse psychological frameworks can strengthen LLM design, enhance alignment, and broaden the practical and ethical impact of modern NLP systems.
As shown in Figure 1, the remainder of this paper illustrates how cognitive, developmental, behavioral, social, psycholinguistic, and personality theories integrate into four key stages of LLM development: preprocessing (Section 2), pre-training (Section 3), post-training (Section 4), and evaluation and application (Section 5). Finally, Section 6 examines three overarching questions: How does current LLM development leverage psychological theories? Which untapped psychological insights could advance LLM development? And what debates loom at the intersection of NLP and psychology?
{forest}
[…]
Figure 1: Our structured survey of how psychological theories apply across the main stages of LLM development. Colors indicate six distinct psychology areas: red for Developmental Psychology; orange for Behavioral Psychology; yellow for Cognitive Psychology; green for Social Psychology; blue for Personality Psychology; purple for Psycholinguistics.
2 Preprocessing
We begin our stage-by-stage analysis of LLM development with preprocessing, the foundation that shapes downstream capabilities. During data collection and processing, psychology provides valuable frameworks into human learning and filtering, underscoring the need for realistic, developmentally informed datasets and effective filtering strategies.
Data Collection  Recent evidence shows that LLMs can align with human brain responses under biologically plausible training conditions (Hosseini et al., 2024), despite LLMs typically requiring orders of magnitude more training data than humans receive. This supports the application of ecological validity (Schmuckler, 2001) that emphasizes real-world data to mimic cognitive development. Drawing from this principle, Jagadish et al. (2024) selects linguistically diverse environments to reflect children’s language acquisition processes, while Feng et al. (2024) utilizes child-directed speech to mirror naturalistic learning contexts. In parallel, developmental theories like incremental numerical understanding (Piaget, 2013) that views numerical concepts as gradually acquired through exposure offer insights into sequential data organization. Applying this framework, Sharma et al. (2024) introduces mathematically coherent numeric anchors to align the collection process.
Data Preprocessing  Data preprocessing inspired by cognitive psychology involves refining data to enhance informational coherence prior to training. Selective attention (Treisman, 1969) – distinct from the attention mechanisms in transformers – prioritizes cognitively salient information while filtering out irrelevant stimuli. Implementing this principle, Nottingham et al. (2024) develops a preprocessing model that identifies and filters irrelevant data. Meanwhile, predictive coding (Rao & Ballard, 1999) proposes anticipatory processing based on prior knowledge, providing a framework for structuring sequential information. Leveraging this insight, Araujo et al. (2021) structures sentences to enable anticipation of subsequent content, improving semantic coherence through expectation-driven processing.
3 Pre-Training
Building on the foundations established during preprocessing, pre-training in LLMs mirrors aspects of human cognitive development, where core linguistic and reasoning abilities emerge through exposure to stimuli. This section explores how psychological insights inform observational learning and knowledge acquisition in LLMs.
Observational Learning and Self-Supervision
Incremental cognitive development (Piaget, 1976), which posits children acquire knowledge through sequential tasks, informs how LLMs can master nuanced concepts with explicit structured exposure. This principle manifests in Schulze Buschoff et al. (2023)’s gradually expanding pre-training tasks and Chen et al. (2024d)’s use of contradictory historical tasks for conceptual restructuring. Additionally, scaffolding theory (Park & Reuter-Lorenz, 2009), which emphasizes gradually challenging interactions, addresses the challenge of maintaining coherent learning trajectories – demonstrated by Borges et al. (2024)’s structured feedback loops and Sonkar et al. (2023)’s dynamic adjustment of task complexity based on performance metrics.
World Knowledge Acquisition
Semantic coherence during pre-training draw insights from top-down and bottom-up perception (Gregory, 1997), which frames cognition as interaction between conceptual frameworks and detailed data. Top-down processing is leveraged to prioritize high-level semantic processing before syntactic details (Rawte et al., 2022) and to generate test case (Zhang et al., 2024b). Meanwhile, Pang et al. (2023) fuses granular bottom-up encoding with top-down corrections for document summarization.
4 Post-Training and Alignment
With foundational knowledge acquired in pre-training, post-training refines LLMs from general proficiency to task-specific, goal-oriented behavior. We explore how psychology guide post-training setups for more context-aware, interpretable, and human-aligned outcomes.
Supervised Fine-Tuning (SFT)  In SFT, works that draw on psychological insights focus on retaining contextual information. Building on working memory theory (Baddeley & Hitch, 1974b) that proposes a short-term system for temporarily holding information, Kang et al. (2024) adds a module into Decision Transformers, enabling LLMs to retain and process short-term information. Beyond retaining, Li et al. (2023) proposes a working memory approach that dynamically balances stored information with provided contexts for SFT.
Reinforcement Learning from Human Feedback (RLHF) A classic behavioral theory, the Operant Conditioning theory posits that behaviors are systematically strengthened or weakened by the consequences (rewards or punishments) that immediately follow them (Thorndike, 1898; Skinner, 1957). The principles of reinforcement learning align closely with this psychological framework, particularly in the post-training phase of LLM development, where RLHF explicitly operationalizes Operant Conditioning theory to align model behaviors with human values and preferences. Through repeated feedback, the model gradually adapts to favor outputs that consistently yield higher reward signals, thereby aligning closer with human preferences – reflecting Thorndike’s Law of Effect, which asserts that behaviors followed by satisfying outcomes are more likely to recur (Lambert et al., 2023). During RLHF, the model generates responses, and a learned reward function $R(x)$ assigns scores to outputs $x$, guiding subsequent policy updates. For instance, Ouyang et al. (2022) train InstructGPT using Proximal Policy Optimization (Schulman et al., 2017), rewarding responses preferred by humans and penalizing less desirable ones. Foundational frameworks (Christiano et al., 2017; Sutton & Barto, 2018; Stiennon et al., 2022) established methods for explicitly translating human judgments into reward signals, operationalizing the insights of Operant Conditioning. More recent work incorporates human cognitive biases (Siththaranjan et al., 2024) and personalizes reward functions for individual values (Poddar et al., 2024). These developments illustrate how Operant Conditioning remains central to aligning LLMs with nuanced human values. While our survey focuses on psychological dimensions, a technical overview of RLHF methods is provided in Appendix B.
5 Evaluation and Application
In the evaluation and application stage of LLM development, psychology offers valuable tools for both assessing and enhancing model behavior. In this section, we review three key areas of challenges where psychology can inform this stage: (1) evaluating emergent capabilities such as reasoning, (2) improving task performance in domains involving human cognition, and (3) designing socially aware, multi-agent systems.
5.1 Benchmarks and Capability Assessment
Evaluating LLMs with psychologically grounded metrics offers a deeper window into their real-world viability. By mapping classic theories onto benchmarks that probe model responses under diverse, human-like scenarios, researchers can move beyond surface-level performance measures, revealing emergent model behavior and illuminating strengths, blind spots, and opportunities to refine LLM training and alignment practices.
5.1.1 Social Reasoning and Intelligence
Social intelligence is vital for LLMs that navigate human contexts, enabling the interpretation of implicit cues, adaptation to social norms, and authentic interaction – defining advanced AI beyond mere text prediction. As LLMs increasingly mediate communication, their grasp of social dynamics becomes pivotal for both efficacy and safety.
Notably, Theory of Mind (ToM) offers a framework for evaluating how individuals understand and attribute mental states – such as beliefs, desires, and intentions – to others (Baron-Cohen et al., 1985). By measuring an LLM’s capacity to represent and reason about beliefs, researchers can assess core social intelligence. Recent benchmarks include ToMBENCH (Chen et al., 2024c), OpenToM (Xu et al., 2024a), HI-TOM (Wu et al., 2023), and FANTOM (Kim et al., 2023), each probing distinct facets of ToM. Extending these efforts to spoken dialogues, Soubki et al. (2024) reveal lingering gaps between LLM and human performance. Surveys (Ma et al., 2023; Sap et al., 2022) then consolidate methods and underscore the challenges of building robust ToM-based evaluations.
Beyond individual cognition, social influence theories, such as Conformity Theories, capture how group pressure shapes individual judgments (Asch, 2016). Recent work tests LLM-based agents’ collaboration and bias dynamics under these principles (Zhang et al., 2024c; Jin et al., 2024b), bridging individual and group-level cognition.
Emotion recognition is another pillar of social intelligence. Ekman’s Basic Emotion Theory (Ekman, 1992) identifies six universal emotions, often used for labeling (Gong et al., 2024), while Dimensional Models like the Circumplex Model conceptualize emotions along valence and arousal (Morrill et al., 2024). LLMs increasingly excel at emotion recognition, benefitting dialogue and sentiment tasks (Zhang et al., 2024e; Wu et al., 2024c; d; Sabour et al., 2024).
These efforts collectively demonstrate both progress and limitations in LLMs’ social cognition, establishing benchmarks against which future developments can be measured.
5.1.2 Language Proficiency
Recent work adopts psycholinguistic assessments, originally designed for humans, to test LLMs’ language proficiency. These experiments probe a wide range of linguistic domains: morphology (Anh et al., 2024), syntax (Liu et al., 2024b; Hale & Stanojević, 2024), phonology (Duan et al., 2025), semantics (Duan et al., 2025; Hayashi, 2025) and their interactions (Miaschi et al., 2024; Zhou et al., 2025).
Although LLMs exhibit comparable performance to human speakers on many psycholinguistic tasks, the underlying processing mechanism they rely on may seem different from humans (Lee et al., 2024). Human language acquisition is often characterized by the Poverty of the Stimulus, i.e. children acquire complex grammar from relatively little input (Chomsky, 1980), whereas LLMs typically require developmentally implausible amounts of linguistic data to learn morphological rules. On the other hand, some evidence suggests that the learning patterns of LLMs mirror aspects of human language acquisition (Liu et al., 2024b).
Several studies have explored the pragmatic abilities of LLMs, motivated by the close link between language and broader cognitive functions in humans. Grice (1975)’s Theory of conversational implicature posits that utterance interpretation depends on both literal content and surrounding context. Researchers (Bender & Koller, 2020; Gubelmann, 2024) have contrasting perspectives on LLMs with respect to the Harnad (1990)’s Symbol Grounding Problem, i.e. linguistic symbols must be grounded in sensorimotor interactions to be meaningful. Failures of LLMs in pragmatic and semantic tasks (Kibria et al., 2024; Zeng et al., 2025), as well as their neuron patterns (Wu et al., 2024b), point to limitations beyond pure linguistic knowledge, which potentially parallel human higher-level cognitive processes.
5.1.3 Memory and Cognitive Evaluation
Assessing memory and cognition is crucial given LLMs’ limited capacity and risk of catastrophic forgetting. Cognitive load is measured by Xu et al. (2024b) on jail-breaking and Zeng et al. (2024b) on memorization patterns. Memory is measured by Li et al. (2023) on parametric knowledge, by Zhang et al. (2024a) on n-back tasks and by Timkey & Linzen (2023) on capacity. Meanwhile, cognitive development and reasoning capabilities have been assessed through cognitive maturity (Laverghetta Jr. & Licato, 2022), subjective similarity (Malloy et al., 2024), reasoning strategies (Mondorf & Plank, 2024; Yuan et al., 2023), decision-making (Ying et al., 2024), and ToM (Jung et al., 2024).
5.1.4 Personality Capability
Personality consistency examines how stably LLMs maintain traits across contexts. Frisch & Giulianelli (2024) show LLMs with asymmetric profiles vary in Big Five traits – Openness, Conscientiousness, Extraversion, Agreeableness, Neuroticism, while Amidei et al. (2025) find language switching alters GPT-4o’s traits on the Eysenck Personality Questionnaire Revised, underscoring challenges in preserving stable traits and reducing context dependence. Parallel research examines how LLMs display and control personality traits. Jiang et al. (2024) show LLMs can express distinct Big Five traits recognized by human evaluators. Mao et al. (2024) introduce PersonalityEdit, revealing difficulties in maintaining consistent alignment for Neuroticism, Extraversion, and Agreeableness, while Hu & Collier (2024) find persona-based prompting improves annotation accuracy.
5.1.5 Bias and Ethics Evaluation
Evaluating biases and ethical risks is crucial for responsible AI that avoids reinforcing harmful social patterns. As LLMs increasingly shape public discourse, thorough assessments are essential to prevent discriminatory outputs and promote equitable benefits across diverse communities. Recent work tests LLMs on gender (Oba et al., 2024; Zhao et al., 2024), broader social biases (Shin et al., 2024; Lee et al., 2023; Nozza et al., 2022), toxic content (Gehman et al., 2020; Luong et al., 2024; Hui et al., 2024), and harmful stereotypes (Shrawgi et al., 2024; Huang & Xiong, 2024; Grigoreva et al., 2024), establishing benchmarks across cultures and languages. Evidence also suggests that LLMs replicate social identity biases, mirroring human tendencies toward ingroup favoritism and outgroup hostility (Hu et al., 2025a; Dong et al., 2024a) – patterns central to social identity theory, which posits that group membership shapes self-concept and intergroup behavior (Tajfel, 1979).
5.2 Task Performance Enhancement
Building on the benchmarks, we review how psychological insights boost LLM task performance. We highlight techniques that enhance complex reasoning and enrich dialogue, illustrating how psychology improves capabilities and alignment across applications.
5.2.1 Reasoning Enhancement
LLMs often struggle with complex reasoning – social inference (Liu et al., 2024a), logical errors (Turpin et al., 2023; McKenna et al., 2023), hallucinations (Huang et al., 2025; Ai et al., 2024a), and multi-step planning (Wang et al., 2024a). Psychological theories of human cognition offer frameworks to address these issues by implementing analogous cognitive mechanisms. For instance, Dual-process theories, a social cognition framework, distinguish between fast (System 1) and slow (System 2) reasoning (Kahneman, 2011), offering a blueprint for LLM improvement. Chain-of-thought prompting (Wei et al., 2022) operationalizes System 2 via intermediate steps, while DynaThink (Pan et al., 2024) dynamically selects rapid or thorough inference. Tree of Thoughts (Yao et al., 2024) further explores multiple reasoning paths concurrently.
Similarly, Self-reflectionintrospection focused on the self-concept (Phillips, 2020) – has guided LLM enhancements in hallucination mitigation (Ji et al., 2023), translation (Chen et al., 2024a; Wang et al., 2024e), question-answering (Li et al., 2024a; Zhang et al., 2024f; Kassner et al., 2023), and math reasoning (Zhang et al., 2024f). Approaches include iterative self-assessment (Ji et al., 2023; Yan et al., 2024), task decomposition (Wang et al., 2024e; Zhang et al., 2024d), reflection-driven self-training (Dou et al., 2024), and confidence-tuned reward functions (Xu et al., 2024c). Additionally, ToM adaptations boost LLMs’ interpersonal reasoning, aiding missing knowledge inference (Bortoletto et al., 2024), common ground alignment (Qiu et al., 2024), and cognitive modeling (Wu et al., 2024a).
Beyond social reasoning, Perception, Attention, and Memory support coherence and retrieval. Kojima et al. (2022) uses “Let’s think step by step” prompts for top-down reasoning. Maharaj et al. (2023); Yu et al. (2022) leverages selective attention to detect hallucinations and extract relation. Zhu et al. (2024) employs recitation for retrieval, and Park & Bak (2024) introduce separate short- and long-term memory modules. Finally, Wang et al. (2024c); Chi et al. (2023) improve complex reasoning via symbolic and adaptive memory structures.
Another challenge is multi-step reasoning with external knowledge, as in retrieval-augmented generation (RAG). Hippocampal indexing theory (Teyler & DiScenna, 1986), viewing the hippocampus as a pointer to neocortical memory, enhances RAG (Gutierrez et al., 2024) and counterfactual reasoning (Miao et al., 2024a). Meanwhile, self-reflection and meta-cognition (Phillips, 2020; Flavell, 1979), supporting iterative introspection, improve retrieval (Asai et al., 2024) and multi-step inference (Zhou et al., 2024).
5.2.2 Dialogue Understanding and Generation
Dialogue understanding is a key area where personality psychology aids trait-based inferences from user interactions. NLP research has explored dynamic ways to measure personality beyond structured tests. The Myers–Briggs Type Indicator (MBTI), a self-report questionnaire that makes pseudo-scientific claims to categorize individuals into 16 distinct personality types, remains popular (Rao et al., 2023; Yang et al., 2023), while PsychoGAT (Yang et al., 2024) gamifies MBTI, and PADO (Yeo et al., 2025) adopts a Big Five-based multi-agent approach. Beyond assessments, traits guide dialogue generation: Huang & Hadfi (2024) show higher agreeability improves negotiation, while Cheng et al. (2023) reveal social and racial biases in persona creation, raising representational concerns.
Dialogue generation research further incorporates personality to improve coherence, empathy, and consistency. Pal et al. (2025); Chen et al. (2025a) leveraged Reddit-based journal entries to model Big Five traits in large-scale dialogue datasets. Other efforts improve persona consistency without referencing explicit psychological theory (Wu et al., 2025b; Takayama et al., 2025). Similarly, personality is used to improve truthfulness, consistency, and context-aware generation in LLMs, as further detailed in Appendix C. These approaches support personality alignment but lack grounding in deeper psychological theory.
5.3 Collaborative and Multi-Agent Frameworks
Beyond task-specific capabilities, the surge in multi-agent LLM frameworks reflects a growing emphasis on collaborative decision-making – where modeling social dynamics is crucial. Social and personality psychology theories offer key insights into designing agent interaction, negotiation, and consensus, guiding more socially intelligent LLM systems.
Social Influence
Persuasion models (Petty & Cacioppo, 2012) illustrate how central or peripheral routes shape attitudes in collaborative settings. Leveraging this, Gollapalli & Ng (2025) merges persuasive dialog acts with reinforcement learning, Furumai et al. (2024) combines LLM strategies and retrieval, Qin et al. (2024); Jin et al. (2024a) emphasize credibility-aware generation, and Zeng et al. (2024a) uncovers LLM vulnerabilities. Multi-agent research simulates personality-driven negotiation (Huang & Hadfi, 2024; Hu et al., 2025b), boosts truthfulness via structured debates (Khan et al., 2024), and curates argument-strength datasets (Rescala et al., 2024).
Social Cognition
ToM complements social influence by enabling agents to grasp others’ mental states. Some efforts integrate belief tracking (Sclar et al., 2023) and coordination (Wang et al., 2022; Sclar et al., 2022), while recent approaches (Wilf et al., 2024; Jung et al., 2024) refine ToM via task decomposition and recursive simulation (Sarangi et al., 2025).
Role-Play and Multi-Agent Simulation
Recent work on persona-driven LLM agents focuses on simulating diverse perspectives, persona alignment, and socially intelligent interactions. Han et al. (2024) introduce PSYDIAL (Big Five-based Extraversion), Castricato et al. (2025) present PERSONA (1,586 synthetic personas), and Wu et al. (2025a) release the RAIDEN Benchmark (40K multi-turn dialogues). Agents also model opinion dynamics (Wang et al., 2025) and evaluate social intelligence (Chen et al., 2024b), with RoleLLM (Wang et al., 2024b), Character100 (Wang et al., 2024d), and persona-aware graph transformers (Mahajan & Shaikh, 2024) further supporting multi-party simulations. Additionally, Kumarage et al. (2025) simulate multi-turn social engineering attacks with LLM agents of varied personality traits, highlighting how psychological profiles shape user vulnerability and the need for more robust defenses against personalized manipulation.
6 Trends and Discussion
6.1 How Does Current LLM Development Harness Psychological Theories?
From the review from Section 2-5, we observe psychological theories have been incorporated into LLM development in stage-specific ways, with uneven coverage across theoretical domains. Figure 1 maps this integration across the development stages.
In early stages, data selection and pretraining, developmental psychology is often referenced. Its emphasis on staged learning and knowledge acquisition aligns with curriculum learning and progressive data exposure, mirroring human developmental trajectories. In post-training, especially RLHF, behavioral psychology ideas are most prominent. Conditioning, reinforcement schedules, and reward design are commonly used to guide model alignment with human preferences. In evaluation and application, theories from social psychology, personality psychology, and psycholinguistics are commonly cited, reflecting a focus on interaction patterns, user modeling, and linguistic variation – areas traditionally explored within these sub-fields. Their prominence in later stages aligns with their emphasis on human-centered communication. Cognitive psychology appears across all stages, particularly in modeling internal mechanisms such as reasoning, memory, and attention. Its breadth makes it a foundational influence, though less visible in agentic interaction settings.
The observed unevenness in integration reflects, perhaps a gap, but more probably a functional alignment – some psychological domains are naturally better suited for certain stages of model development. Meanwhile, these trends expose under-explored opportunities, which motivate the research questions that follow.
6.2 What Untapped Psychological Insights Could Advance LLM Development?
Although psychological theory is increasingly applied in LLM research, its use remains simplified and uneven. As shown in Tables A, A, and A, many theories are under-utilized despite their potential to improve model behavior and interpretability. Below, we outline theories in four key areas that deserve greater attention in future LLM research.
Social psychology remains underutilized in areas like group dynamics and self and identity, limiting personalization, adaptability, and inclusivity. Prompting LLMs to adopt specific social identities can reduce bias (Dong et al., 2024a) and mirror human-like ingroup favoritism (Hu et al., 2025a). Incorporating social identity frameworks could enhance user alignment in identity-sensitive contexts (Chen et al., 2020). Likewise, while bias detection is common, classic social influence theories (e.g., conformity, obedience) and attitude change theories (e.g., balance theory, cognitive dissonance) are rarely applied to interaction dynamics or bias mitigation, despite their relevance to ethical and socially adaptive behavior. Additionally, malicious actors leveraging social influence can severely undermine trust in digital spaces (Zeng et al., 2024a; Liu et al., 2025; Ai et al., 2024b), highlighting the potential of constructs like inoculation theory to proactively guard against manipulative strategies.
Behavioral psychology inspires RLHF, yet key concepts like partial reinforcement, which improves behavior persistence (Ferster, 1966; Jensen, 1961), and shaping, which supports gradual learning through successive approximations (Love et al., 2009), are overlooked. Current RLHF relies on uniform rewards, yet behavioral theory warns that flawed rewards can lead to reward hacking. Adding reward variability may reduce premature convergence and improve alignment with human intent (Dayan & Daw, 2008; Amodei et al., 2016).
Personality Psychology use focuses on Trait Theory, overlooking developmental theories that explain how individual traits emerge, evolve, and adapt across contexts. These developmental models could enable more coherent and interpretable personality representations, offering a deeper alternative to static prompt-based personas.
Cognitive psychology remains underused, particularly Schema Theory, which holds that humans store knowledge as dynamic, structured representations formed through repeated experience (Anderson & Pearson, 1984), guiding inference, memory, and learning. Recent work explores schema-inspired methods for compressing user histories and modeling knowledge activation cycles (Panagoulias et al., 2024; Xia et al., 2024), though these remain peripheral. Further integration may improve long-term context handling and generalization.
6.3 What Debates Loom at the NLP–Psychology Intersection, and Where Next?
A recurring question is whether human psychology can be directly mapped to LLMs without distortion (Löhn et al., 2024). Below, we highlight key controversies at this boundary; see Appendix D for an extended discussion. These challenges motivate new recommendations and highlight open directions for cross-disciplinary exploration.
Terminology Mismatches
A core tension is the mismatch between psychological terminology and their NLP usage. For example, attention in psychology means selective mental focus, but in transformers it is a token weighting mechanism without cognitive awareness (Lindsay, 2020), leading to misleading attributions of intentionality. Similarly, memory in psychology entails structured encoding and recall, whereas in LLMs it typically refers to context windows or parameters. Such anthropomorphic language is increasingly prevalent and shapes public and scholarly assumptions about LLMs, as recent studies show rising human-like descriptors (Ibrahim & Cheng, 2025). This calls for disentangling metaphor from mechanism through a precise cross-disciplinary lexicon, preventing both oversimplification and over-anthropomorphization – an underexplored but crucial research challenge.
Theoretical Discrepancies in Use of Psychology
Beyond terminology, deeper theoretical mismatches arise when the NLP community adopts outdated or disputed concepts from psychology. For instance, predictive coding (Rao & Ballard, 1999) is used to analogize LLMs’ next-token prediction, although current research emphasizes hierarchical, multi-scale brain mechanisms (Antonello & Huth, 2024; Caucheteux et al., 2023). Likewise, folk-psychological typologies like MBTI persist in LLM applications despite its criticized validity and reliability (Pittenger, 1993; McCrae & Costa Jr, 1989).
Working memory (Baddeley & Hitch, 1974a) illustrates another gap: LLM ’memory’ modules (Kang et al., 2024; Li et al., 2023) do not replicate human constraints, prompting questions about whether AI should emulate human cognitive limits or exceed them for performance gains. Behavioral psychology faces similar critiques (Miller, 2003; Flavell et al., 2022), as RLHF often focuses on reward optimization (Ouyang et al., 2022; Rafailov et al., 2023; Ramesh et al., 2024), neglecting internal states and risking reward hacking (Skalse et al., 2022; Krakovna, 2020). Broader debates remain over whether LLMs truly “understand” language or function as “stochastic parrots” (Ambridge & Blything, 2024; Park et al., 2024).
In response, we recommend refining how psychological theories are mapped into computational models, replacing outdated constructs with supported frameworks, exploring whether human-like constraints aid interpretability, and designing evaluations that track both outputs and internal states. Sustained collaboration between computational and psychological sciences is essential for robust and theory-aligned LLMs.
Evaluation and Validity Debates
Another major debate is how we evaluate LLM “psychological” abilities – whether current tests really measure what they claim. For instance, GPT-4 solves around 75% of false-belief tasks, matching a 6-year-old’s performance (Kosinski, 2024; Strachan et al., 2024); some see emergent ToM-like reasoning (Kosinski, 2024), but others argue it may be pattern matching (Strachan et al., 2024), noting that minor prompt changes can derail results (Shapira et al., 2024). This calls for more theory-grounded evaluation and clearer definitions.
A parallel controversy involves personality: some studies find stable simulated traits (Sorokovikova et al., 2024; Huang et al., 2024), while others reveal variability under different prompt conditions (Gupta et al., 2024; Shu et al., 2024), raising questions about inherent vs. mimicked personas (Tseng et al., 2024). These debates underscore the need for a systematic, theory-driven framework beyond surface metrics, guiding more faithful replication of human cognition and behavior in LLMs.
7 Conclusions
We systematically review how psychology can ground LLM innovation in both past and future across key subfields: cognitive, developmental, behavioral, social, personality, and psycholinguistic. We examine how psychological theories inform each stage of LLM development, revealing both meaningful connections across domains and critical points of tension, which we explore through discussion to help bridge interdisciplinary gaps. We hope this survey sparks reflection, and inspires future work to continue integrating psychological perspectives into NLP in meaningful and impactful ways.
Limitations
Our survey primarily focuses on literature within NLP, particularly in how personality is modeled, evaluated, and leveraged in LLMs. As a result, we do not extensively cover research from psychology and cognitive sciences that might offer deeper theoretical insights into human-like behaviors in AI. This limitation may exclude valuable methodologies or perspectives that could enhance personality evaluation frameworks for LLMs. We encourage future surveys to integrate findings from psychology and linguistics to bridge theoretical foundations with computational approaches, fostering a more comprehensive understanding of personality in AI systems.
While our survey advocates for a deeper integration of psychology into LLM design, we also caution against the ethical risks posed by overuse or misapplication of psychological principles. A concrete example is operant conditioning (Skinner, 1957), which describes how behavior can be shaped by consequences. Applied to LLMs – for instance, through timely, gratifying feedback to reinforce engagement – these mechanisms can be beneficial in contexts like language learning or motivation. However, reinforcement schedules such as variable ratio or interval rewards may unintentionally condition users to engage compulsively, raising the risk of manipulative design. This presents a key ethical limitation: distinguishing between genuinely supportive interactions and those that encourage excessive use is inherently difficult. To address this, we emphasize the need for transparent disclosure of reinforcement mechanisms and the establishment of clear ethical guidelines by professional communities. These safeguards are essential to ensure that psychological insights enhance user well-being without enabling exploitative practices.
Appendix
Appendix A Psychology Theories
{NiceTabular}
C65pt—c—C80pt—C250pt—c Psych Area Sub Area Theory Definition Explored
[…]
Scaffolding Theory Learning is enhanced through gradually challenging interactions with appropriate guidance (Park & Reuter-Lorenz, 2009)
Incremental Numerical Understanding Numerical concepts are gradually acquired through structured exposure and experience (Piaget, 2013)
Zone of Proximal Development Optimal learning occurs in the gap between what a learner can do independently and with assistance (Wertsch, 1988)
[…]
Ecological Validity Emphasizes real-world data and environments to mimic natural cognitive development (Schmuckler, 2001)
[…]
Top-down and Bottom-up Processing Distinguishes between concept-driven (top-down) and data-driven (bottom-up) perceptual processing (Gregory, 1997)
Predictive Coding Anticipatory processing based on prior knowledge and prediction of expected inputs (Rao & Ballard, 1999)
[…]
Long-term Memory System for storing information over extended periods through semantic organization (Tulving et al., 1972)
Hippocampal Indexing Theory Views the hippocampus as a pointer to neocortical memory representations (Teyler & DiScenna, 1986)
[…]
Theory of Mind The ability to attribute mental states to oneself and others and understand others may have different beliefs (Baron-Cohen et al., 1985)
Schema Theory Knowledge is organized into interconnected patterns that guide processing and interpretation of new information (Anderson & Pearson, 1984)
[…]
Operant Conditioning Behavior is strengthened or weakened by consequences such as rewards or punishments (Skinner, 1957; 1938)
Thorndike’s Law of Effect Behaviors followed by satisfying outcomes are more likely to be repeated in the future (Thorndike, 1927)
Premack Principle A preferred activity can reinforce a less preferred one if access is contingent (Premack, 1959)
Table 1: Representative developmental, cognitive, and behavioral psychology theories by sub-area. In the “Explored” column,
indicates multiple surveyed works,
indicates fewer than three, and
indicates that none emerged in our survey (i.e., not yet substantially explored).
{NiceTabular}
C65pt—c—C80pt—C250pt—c Psych Area Sub Area Theory Definition Explored
[…]
Dual-Process Theory Differentiates between fast, intuitive (System 1) and slow, deliberate (System 2) reasoning (Kahneman, 2011)
Theory of Mind (ToM) How individuals understand and attribute mental states to others (Baron-Cohen et al., 1985)
[…]
4-1Social Influence Social Impact Theory The magnitude of social influence depends on the strength, immediacy, and number of sources (Latané, 1981)
Conformity Theories Explore how group pressure can alter individual judgments (Asch, 2016)
Obedience Theories Demonstrate how authority influences behavior, highlighting conditions under which individuals comply (Milgram, 1963)
Persuasion Models Explain how messages processed via central or peripheral routes can lead to attitude change (Petty & Cacioppo, 2012)
[…]
2-1Group Dynamics Groupthink Examines how the desire for conformity and group cohesion can lead to flawed decision-making and suppression of dissenting opinions (Janis, 1972)
Social Facilitation and Social Loafing Investigates how the presence of others can enhance performance on simple tasks or reduce effort in collective work (Zajonc, 1965; Latané et al., 1979)
[…]
4-1Attitude Change Cognitive Dissonance Theory Explains how inconsistencies between beliefs or behaviors create discomfort, prompting attitude change to restore consistency (Morvan & O’Connor, 2017)
Elaboration Likelihood Model (ELM) Proposes that persuasion occurs via a central route (deliberate processing) or a peripheral route (heuristic processing), depending on the recipient’s motivation and capacity (Petty & Cacioppo, 2012)
Balance Theory Suggests that individuals strive for consistency among their attitudes and relationships, adjusting beliefs to maintain cognitive harmony (Heider, 1946)
Inoculation Theory Posits that exposure to weak counterarguments can strengthen resistance to persuasion by preemptively activating defensive mechanisms (McGuire, 1964)
[…]
5-1Self and Identity Self-Reflection Defines the process of introspection, with attention placed on the self-concept (Phillips, 2020)
Self-Perception Theory Explains how individuals infer their internal states by observing their own behavior (Bem, 1972)
Social Identity Theory Posits that group membership shapes self-concept and influences intergroup behavior (Tajfel, 1979)
Self-Categorization Theory Expands on social identity theory, describing how individuals classify themselves and others into social groups, shaping social norms (Maines, 1989)
Self-Affirmation Theory Suggests that individuals are motivated to maintain their self-integrity when faced with threats to their self-concept (Steele, 1988)
Table 2: Representative social psychology theories by sub-area. In the “Explored” column,
indicates multiple surveyed works,
indicates fewer than three, and
indicates that none emerged in our survey (i.e., not yet substantially explored).
{NiceTabular}
C65pt—c—C80pt—C250pt—c Psych Area Sub Area Theory Definition Explored
[…]
Myers-Briggs Type Indicator (MBTI) Classifies individuals into 16 personality types based on four dichotomies (e.g., Introversion vs. Extraversion) (Myers & Myers, 1995). While widely used, MBTI has been criticized for lacking empirical validity, reliability, and independence between its categories. (Pittenger, 1993)
Eysenck Personality Questionnaire-Revised (EPQR-A) Contains a 24-item personality test that measures extraversion, neuroticism, psychoticism, and social desirability. (Eysenck & Eysenck, 1984)
[…]
5-1Personality Theories Humanistic Theory Emphasizes free will, personal growth, and self-actualization. This perspective focuses on individuals’ subjective experiences and their drive to achieve their full potential. (Stefaroi, 2015)
Psychoanalytic Theory Originating from Freud, this theory conceptualizes personality as the dynamic interplay between the id, ego, and superego, with unconscious processes playing a central role in shaping behavior. (Scharff et al., 2013)
Behaviorist Theory Views personality as a set of learned responses shaped by environmental reinforcements and punishments. This perspective, pioneered by figures like Skinner and Watson, rejects internal mental states in favor of observable behaviors. (Pierce & Cheney, 2008)
Social Cognitive Theory Highlights the role of cognitive processes in personality, emphasizing how expectations, beliefs, and observational learning shape behavior. (Spielman et al., 2024)
Trait Theory Focuses on identifying and measuring stable personality traits that influence behavior across different contexts. (Cartwright, 1979)
[…]
Usage-Based Theory Emphasizes the role of social interaction and cognitive processes in language learning, rather than innate universal grammatical structures (Tomasello, 2005)
[…]
Constraint-Based Models Language processing is an interactive, probabilistic process where multiple sources of information simultaneously contribute to understanding, rather than following a strict, sequential parsing approach (MacDonald et al., 1994)
Good-Enough Processing Proposes that humans comprehend language through approximate, semantically-focused representations that capture the core meaning rather than constructing syntactically perfect linguistic interpretations (Ferreira & Patson, 2007)
Construction-Integration Model Describes text comprehension as a two-stage process where readers first generate multiple, loosely connected propositions and then systematically filter and integrate them into a coherent, meaningful understanding. (Kintsch, 1988)
[…]
Interactive Two-Step Model An interactive, probabilistic process of lexical selection and phonological encoding, where multiple linguistic levels simultaneously influence each other during speech generation (Goldrick & Rapp, 2007)
Table 3: Representative personality psychology and psycholinguistics theories by sub-area. In the “Explored” column,
indicates multiple surveyed works,
indicates fewer than three, and
indicates that none emerged in our survey (i.e., not yet substantially explored).
In Tables A, A, and A, we summarize representative theories in each psychological domain and note whether they have been leveraged in LLM development.
Appendix B Extended Discussion on Reinforement Learning from Human Feedback (RLHF)
B.1 Operant Conditioning in RLHF
During RLHF fine-tuning, the model (agent) generates responses while a learned reward function $R(x)$, often a neural network trained on preference data, assigns scores to candidate outputs $x$. These scores proxy for human judgment and guide policy updates to reinforce higher-reward behaviors. For instance, Ouyang et al. (2022) trains InstructGPT via Proximal Policy Optimization (Schulman et al., 2017): responses deemed more helpful or accurate by human evaluators receive greater reward, whereas undesirable or incorrect outputs face penalization. Unlike purely exploration-based RL methods, this arrangement leverages human insight to provide a more precise learning signal; however, success relies on careful and consistent reward modeling that captures subtle human values.
B.2 Modeling Human Preferences as a Reward Function
Although extensive work has been conducted in RLHF, here we primarily highlight recent approaches or methodologies explicitly grounded in psychological theories. Building robust reward functions from heterogeneous or ambiguous feedback remains a core challenge in RLHF. Early foundational frameworks (Christiano et al., 2017; Stiennon et al., 2022) laid essential groundwork for converting human judgments into usable reward signals, drawing implicitly from principles of Operant Conditioning Theory. More recent advancements explicitly target improvements in stability, scalability, and fairness, addressing issues arising from the inherent variability and complexity of human preferences.
Rafailov et al. (2023) introduced Direct Preference Optimization (DPO), simplifying preference integration by directly optimizing the policy through a closed-form solution, thus removing the need for explicit intermediate reward modeling. Extending these efforts toward equitable alignment, Ramesh et al. (2024) proposed Group Robust Preference Optimization (GRPO), ensuring robustly aligned outcomes across diverse demographic groups, addressing biases commonly observed in human-driven reward processes.
Further refinements emphasize enhancing alignment accuracy through psychological considerations. For instance, Contrastive Preference Learning (Hejna et al., 2024) utilizes regret-based losses inspired by behavioral economics, facilitating stable off-policy learning without conventional RL techniques. Distributional Preference Learning (Siththaranjan et al., 2024) aligns reward modeling more closely with human cognitive patterns by capturing human values as probability distributions rather than point estimates. Variational Preference Learning (VPL) (Poddar et al., 2024) further integrates psychological realism, introducing latent-variable modeling to personalize RLHF, reflecting variability in individual user preferences rather than imposing a universal reward structure.
These advancements collectively illustrate how psychological theory, particularly Operant Conditioning Theory, continues to shape and inspire sophisticated techniques for reliably aligning LLM behavior with nuanced human values.
B.3 Reinforcement Schedules and Feedback Frequency
In early RLHF, feedback is typically sparse — a single scalar reward per output — which causes a credit assignment problem: the model can’t tell which parts of the output led to the reward. This is similar to delayed feedback in animal learning, which slows progress. Psychology shows that immediate and frequent reinforcement improves learning. Similarly, recent RLHF methods provide dense, token-level feedback (e.g., from a critic model), which improves sample efficiency and training stability. To address this, Cao et al. (2024) propose LLM self-critique, a method that uses a secondary model to provide dense, token-level feedback during generation. This simulates a continuous reinforcement schedule, analogous to real-time feedback in behavioral training, and leads to more stable and efficient learning. Another factor is how often feedback is given: continuous vs. partial reinforcement. While human feedback is often sparse due to cost, using AI feedback models (like RLAIF, will discuss later) allows for more frequent feedback. Even with limited human scores, techniques like credit assignment can distribute reward across the output.
B.4 Reward Prediction Errors as a Learning Driver
At the heart of reinforcement learning lies the concept of reward prediction error (RPE), which arises when there is a discrepancy between an agent’s expected reward and the reward it actually receives, prompting adjustments and driving learning (Sutton & Barto, 2018). This mechanism closely parallels dopaminergic signaling in animal brains, where dopamine neurons respond strongly to unexpected rewards or punishments, effectively reinforcing behaviors associated with positive surprises or reducing those linked to disappointments (Schultz, 1998). In RLHF, reward prediction errors similarly guide model updates; each model output receives a score from a reward model trained on human preferences, and deviations between these scores and the model’s predicted rewards are used to adjust behavior. However, simplistic or flawed reward models can lead to ”reward hacking,” where the model exploits blind spots in the reward function rather than genuinely aligning with human values (Amodei et al., 2016). Introducing variability in reward signals can encourage exploration and mitigate premature convergence on suboptimal strategies (Dayan & Daw, 2008). To address reward hacking and reward-model inconsistencies, recent approaches have formulated RLHF as a constrained Markov decision process with dynamic weighting (Moskovitz et al., 2024), introduced information-theoretic regularization techniques (InfoRM) (Miao et al., 2024b), and proposed methods such as ConvexDA and reward fusion to stabilize and enhance reward-model consistency (Shen et al., 2024).
B.5 Implications for Bias, Alignment, and Reward Modeling
Employing these behavioral principles may improve how well RLHF handles biases and achieves robust alignment. For instance, diverse trainers and variable scenarios can prevent conditioning bias, where the model overfits to a narrow segment of human preferences (Sheng et al., 2019). Moreover, shaping and multi-dimensional reward functions can address multiple alignment goals simultaneously (e.g., factual accuracy and polite style), limiting reward hacking.
At the same time, grounding RLHF in behavioral theory highlights persistent pitfalls. Models still lack an intrinsic understanding of human values, and an imprecise reward signal can reinforce superficial behaviors. To mitigate these risks, a cycle of model auditing, reward model refinement, and re-training can mirror how animal trainers continually adjust reinforcement to avoid unwanted side effects.
Appendix C Persona-Inspired Dialogue Generation
Personality has also inspired improvements truthfulness, response grounding, and broader alignments. Zhang et al. (2024d) introduced Self-Contrast to enhance internal consistency, and Joshi et al. (2024) proposed the Persona Hypothesis, linking truthfulness to pretraining structure. Kim et al. (2024) introduced PANDA to reduce persona overuse in dialogue. Zhang et al. (2024d) introduced a reflection-based technique to reduce internal inconsistencies. Joshi et al. (2024) proposed the Persona Hypothesis, arguing that LLMs encode truthful and untruthful personas from their training distribution. Kim et al. (2024) addressed the overuse of persona cues to improve contextual appropriateness. Persona-guided generation has been applied to emotionally supportive role-play settings (Ye et al., 2025; Chen et al., 2025b).
Appendix D Extended Discussion on Debates over NLP-Psychology Intersection
A recurring theme is whether human psychology can be naively mapped onto LLM behavior without distortion (Löhn et al., 2024). Therefore, in this section, we discuss several major points of contention at this interdisciplinary boundary. These issues motivate a set of recommendations and highlight open directions for future cross-disciplinary research.
Terminology Mismatches
One key issue is the mismatch in terminology and the anthropomorphization of technical concepts. Terms like attention, memory, and “understanding” have specific meanings in psychology that differ from their usage in NLP. For instance, attention in psychology refers to selective mental focus and executive control, whereas in transformers models, it is a mathematical mechanism for weighting tokens – without cognitive awareness (Lindsay, 2020). This divergence can lead to misleading interpretations, such as assuming models exhibit intentional focus when they merely perform matrix operations. Similar misalignments exist for terms like memory (which in psychology implies a structured encoding and recall process, versus an LLM’s context window or weight parameters) and expressions such as “knows” or “thinks.”
Such anthropomorphic language is increasingly prevalent and shapes public and scholarly assumptions about LLMs. Recent analyses have found a growing prevalence of human-like descriptors for LLM behavior, raising calls to carefully disentangle metaphor from mechanism (Ibrahim & Cheng, 2025). An open research direction is developing a more precise cross-disciplinary lexicon: how can we describe model behaviors in ways that neither oversimplify the psychology nor over-anthropomorphize the engineering? Improving interdisciplinary communication by explicitly defining terms and drawing careful analogies remains an important but under-addressed challenge.
Theoretical Discrepancies in Use of Psychology
Beyond terminology, discrepancies arise in the adoption of psychological theories within NLP research. Sometimes, NLP integrates concepts from psychology that are outdated or contested in their original fields. For instance, predictive coding, which proposes that the brain continuously anticipates sensory input and updates via prediction errors (Rao & Ballard, 1999), is often used as a metaphor for LLMs’ next-token prediction. However, contemporary studies emphasize that brain prediction operates across hierarchical and multi-scale structures (Antonello & Huth, 2024; Caucheteux et al., 2023), cautioning against simplistic analogies that risk misrepresenting the theory.
Another example is the lingering use of folk-psychological typologies like the MBTI in some LLM studies. Despite its cultural popularity, MBTI has faced substantial criticism for poor validity and reliability (Pittenger, 1993). It classifies personality into 16 types based on Jungian dichotomies; however, research indicates these categories lack stability and predictive power regarding behavior (McCrae & Costa Jr, 1989). Nonetheless, the ease of obtaining of MBTI-labeled data has led some NLP studies to treat these categories as definitive, highlighting a theoretical lag where NLP adopts psychological models that mainstream psychology has largely moved beyond.
Working memory presents another gap. While cognitive psychology and neuroscience characterize it by limited capacity and active attention control (Baddeley & Hitch, 1974a), LLM approximations – such as short-term retention modules (Kang et al., 2024) or memory mechanisms for external context (Li et al., 2023) – do not replicate these constraints. This raises questions: Should AI systems emulate human cognitive limitations to achieve more human-like reasoning, or should they leverage their capacity to surpass such constraints? If certain human limitations, like bounded memory, lead to desirable properties such as better interpretability or reduced distractions, might it be useful to impose similar limits on AI? These questions remain largely open.
Finally, a related debate concerns behavioral psychology. The field has been critiqued for ignoring cognitive processes (Miller, 2003) and internal mental states (Flavell et al., 2022) that drive the observed behaviors, limiting its explanatory power. With the critiques remaining, the superficial application of behavioral psychology is also evident in LLM research. For instance, RLHF draws from operant conditioning but largely focuses on optimizing rewards (Ouyang et al., 2022; Rafailov et al., 2023; Ramesh et al., 2024), often neglecting internal model states. Consequently, a flip-side of such optimization is reward hacking (Skalse et al., 2022), where models exploit shortcuts without meeting true objectives – mirroring human behavior under evaluative pressure (Krakovna, 2020). Deeper integration of cognitive psychology is needed to address these limitations in LLM design.
The debate over whether LLMs possess a true understanding of language or merely function as ”stochastic parrots” (Bender et al., 2021) remains ongoing. Linguists have largely been skeptical (Ambridge & Blything, 2024), arguing that language ability is inherently abstract and complex, extending beyond mere statistical pattern recognition. Park et al. (2024) connection between mathematical reasoning and high-level linguistic comprehension.
Evaluation and Validity Debates
Anoter central debate concerns how we evaluate LLMs on purportedly “psychological” abilities – and whether current tests measure what we assume. For example, advanced LLMs like GPT-4 perform well on traditional ToM tasks, solving around 75% of false-belief scenarios, comparable to a 6-year-old child (Kosinski, 2024; Strachan et al., 2024). Some interpret this as emergent ToM-like reasoning (Kosinski, 2024), but others caution that high performance may reflect surface-level pattern matching rather than genuine mental-state attribution. Researchers emphasize that correct answers do not imply mentalizing ability (Strachan et al., 2024), and minor prompt changes can significantly impair model performance (Shapira et al., 2024). This underscores the need for more rigorous, theory-grounded evaluations and clearer cross-disciplinary definitions.
A similar controversy surrounds personality modeling. Some studies suggest LLMs exhibit stable simulated personality traits (Sorokovikova et al., 2024; Huang et al., 2024), enabling consistent persona simulation across prompts. However, others show that LLM responses vary with prompt framing and response order, undermining test reliability (Gupta et al., 2024; Shu et al., 2024). Tseng et al. (2024) distinguish between role-playing (adopting assigned traits) and personalization (adapting to users), raising a fundamental question: do LLMs have inherent personalities, or merely mimic behavior? While LLMs can simulate personality, inconsistent assessments cast doubt on whether such traits are emergent or engineered – an open direction for future work.
In summary, these debates highlight the need for a systematic, theory-driven framework that goes beyond superficial performance metrics, thereby enhancing model interpretability and guiding the development of LLMs to more faithfully replicate the complexities of human cognition and behavior.