### v2

# A Review of Incorporating Psychological Theories in LLMs

Abstract
Psychological insights have long shaped pivotal NLP breakthroughs, from attention mechanisms to reinforcement learning and social modeling. As Large Language Models (LLMs) develop, there is a rising consensus that psychology is essential for capturing human-like cognition, behavior, and interaction. This paper reviews how psychological theories can inform and enhance stages of LLM development. Our review integrates insights from six subfields of psychology, including cognitive, developmental, behavioral, social, personality psychology, and psycholinguistics. With stage-wise analysis, we highlight current trends and gaps in how psychological theories are applied. By examining both cross-domain connections and points of tension, we aim to bridge disciplinary divides and promote more thoughtful integration of psychology into NLP research.
A Review of Incorporating Psychological Theories in LLMs
Zizhou Liu∗♠,     Ziwei Gong∗♠,     Lin Ai∗♠,     Zheng Hui∗†,     Run Chen∗♠, Colin Wayne Leach,     Michelle R. Greene,     Julia Hirschberg Equal contributions. Department of Computer Science, Columbia University, Department of Psychology, Barnard College, The Language Technology Lab, University of Cambridge {sara.ziweigong, lin.ai, runchen,julia}@cs.columbia.edu {cleach, mgreene}@barnard.edu,{zl2889, zh2483}@columbia.edu
1 Introduction
As Large Language Models (LLMs) grow in scale and complexity, the Natural Language Processing (NLP) community increasingly sees psychology as key to capturing human-like cognition, behavior, and interaction (Qu et al., 2024; Lewis, 2025). Psychology, grounded in empirically validated and computationally adaptable frameworks (Sartori and Orrù, 2023; Ong, 2024), offer principled tools for addressing persistent LLM challenges such as reasoning fidelity, context retention, and user interaction. Early intersections of psychology and NLP predate modern LLMs. Classic systems such as ELIZA Weizenbaum (1966) framed conversations around simple pattern matching, while early cognitive and psycholinguistic work influenced representation and processing assumptions. Today’s LLMs revive these threads at a new scale, but psychology is best viewed as one productive lens among others, complementary to linguistics, HCI, and sociology, rather than an inevitable or singular foundation. Reflecting these strengths, psychological insights have driven NLP advances, including the cognitive inspirations of attention mechanisms, formative reinforcement learning approaches, and social modeling for agents.
Despite extensive multidisciplinary efforts, a holistic review systematically integrating psychology across the LLM lifecycle remains missing. Most surveys and position papers remain fragmented, typically falling into three broad categories: (1) Some investigate how LLMs can empower traditional psychology or cognitive science research, for instance by modeling human reasoning and behavior at scale (Abdurahman et al., 2024a; Demszky et al., 2023; Ong, 2024; Ke et al., 2024). (2) Others approach LLMs as subjects of psychological analysis, aiming to adapt or extend psychological theory, such as personality or cognition frameworks, to interpret and evaluate model behavior (Brito et al., 2025; Li et al., 2024b; Hagendorff et al., 2023; tse Huang et al., 2024; Pellert et al., 2024). (3) Finally, a third group leverages a single or limited set of psychological constructs to enhance model alignment or multi-agent frameworks – improving system reliability, social interaction, and trustworthiness (Dong et al., 2024b; Liu et al., 2023). This includes research on social influence for AI safety (Zeng et al., 2024a), moral reasoning in legal tasks (Almeida et al., 2024), and partial integrations of social or developmental psychology (Sartori and Orrù, 2023; Serapio-García et al., 2025; Zhang et al., 2024c). However, no existing work provides a unified map of how diverse psychological sub-areas can be harnessed, from data through application. Our survey fills this gap by offering a stage-wise view of how psychology can strengthen LLM capabilities and alignment across the entire lifecycle.
To address this gap, we present a structured review that situates psychological theories from six major areas across the entire LLM development pipeline. The contributions of our survey [^1] We survey 227 papers from major *CL venues, plus COLING, NeurIPS, ICML, ICLR, and influential arXiv preprints. Appendix LABEL:appendix:keywords details paper selection process. are twofold: (1) We systematically review psychological theories applied in key stages of LLM development, identifying gaps and inconsistencies. (2) We highlight under-explored concepts alongside critical issues and debates at the intersection of psychology and NLP. Collectively, these contributions demonstrate how integrating diverse psychological frameworks can strengthen LLM design, enhance alignment, and broaden the practical and ethical impact of modern NLP systems.
As shown in Figure 1, the remainder of this paper illustrates how cognitive, developmental, behavioral, social, psycholinguistic, and personality theories integrate into four key stages of LLM development: preprocessing (Section 1), pre-training (Section 3), post-training (Section 4), and evaluation and application (Section 5). Finally, Section 6 discusses three central questions: (i) How does current LLM development leverage psychological theories? (ii) Which untapped psychological insights could advance LLM development? (iii) What debates loom at the intersection of NLP and psychology?
2 Preprocessing
{forest}
Figure 1: Our structured survey of how psychological theories apply across the main stages of LLM development. Colors indicate six distinct psychology areas: red for Developmental Psychology; orange for Behavioral Psychology; yellow for Cognitive Psychology; green for Social Psychology; blue for Personality Psychology; purple for Psycholinguistics.
We begin the stage-by-stage analysis of LLM development with preprocessing, the foundation that shapes downstream capabilities. Psychology provides valuable frameworks for understanding how humans acquire and filter information, underscoring the need for realistic, developmentally informed datasets and effective filtering strategies in data collection and processing.
Data Construction  Recent evidence shows that LLMs can align with human brain responses under biologically plausible training conditions (Hosseini et al., 2024), despite LLMs typically requiring orders of magnitude more training data than human. This supports the application of ecological validity (Schmuckler, 2001) that emphasizes real-world data to mimic cognitive development. To reflect children’s language acquisition processes, Jagadish et al. (2024) selects linguistically diverse environments, Feng et al. (2024) utilizes child-directed speech, while Nikolaus et al. (2022b) collects child cartoon. In parallel, incremental numerical understanding (Piaget, 1953) that views numerical concepts as gradually acquired through exposure is applied to sequential data collection with mathematically coherent numeric anchors (Sharma et al., 2024). Lastly, Reuben et al. (2025) provides a systematic framework to reformulate psychological questionnaires for LLM assessment.
Data Preprocessing  Data preprocessing inspired by cognitive psychology involves refining data to enhance informational coherence prior to training. Selective attention (Treisman, 1969), prioritizing salient information while filtering out irrelevant stimuli, was implemented to develop a preprocessing model that filters irrelevant data (Nottingham et al., 2024). Meanwhile, predictive coding proposing anticipatory processing based on prior knowledge (Rao and Ballard, 1999), was leveraged by Araujo et al. (2021) to enable anticipation of subsequent content, improving semantic coherence through expectation-driven processing. Lastly, drawing insights from knowledge acquisition of children, Ficarra et al. (2025) redefines lexical knowledge in data to capture distributional information based on target word.
3 Pre-Training
Building on the foundations established during preprocessing, pre-training mirrors human cognitive development, where linguistic and reasoning abilities emerge through exposure to stimuli. This section explores how psychology inform observational learning and knowledge acquisition in LLMs.
Observational Learning Incremental cognitive development (Piaget, 1976), which posits children acquire knowledge through sequential tasks, informs how LLMs can master nuanced concepts with explicit structured exposure. It is used as a motivating analogy for curriculum and sequence effects (Bengio et al., 2009) (i.e., how staged exposure can scaffold capability growth) rather than to claim a theoretical identity with next-token prediction. This principle manifests in Schulze Buschoff et al. (2023)’s gradually expanding pre-training tasks, Chen et al. (2024d)’s contradictory historical tasks and Ma et al. (2025c)’s trial-and-demonstration framework. In contrast, self-supervised objectives align more closely with use-based views of meaning, resonating with Wittgenstein’s claim that meaning emerges through use (Strawson, 1954). Additionally, scaffolding theory (Park and Reuter-Lorenz, 2009), which emphasizes gradually challenging interactions, informs maintaining coherent learning trajectories through Borges et al. (2024)’s structured feedback loops and Sonkar et al. (2023)’s dynamic task complexity.
Knowledge Acquisition Semantic coherence during pre-training draw insights from top-down and bottom-up perception (Gregory, 1997), which frames cognition as interaction between conceptual frameworks and detailed data. Top-down processing is leveraged to prioritize semantic processing before syntactic details (Rawte et al., 2022) and to generate test cases (Zhang et al., 2024b). Meanwhile, to enhance perception modeling, Pang et al. (2023) fuses bottom-up encoding with top-down corrections, and Nikolaus and Fourtassi (2021) models production-based learning. Introducing working memory theory (Baddeley and Hitch, 1974) that proposes a short-term system for temporarily holding information, Mita et al. (2025) simulates critical period dynamics with growing memory capacity to enhance performance.
4 Post-Training and Alignment
With foundational knowledge acquired in pre-training, post-training refines LLMs from general proficiency to task-specific behavior. We explore how psychology guide post-training for context-aware, interpretable, and human-aligned outcomes.
Supervised Fine-Tuning (SFT)  In SFT, works that draw on psychological insights focus on retaining and learn contextual information. Building on working memory theory, Kang et al. (2024) adds a working memory module to retain short-term information, while Li et al. (2023) dynamically balances memory with contexts to improve robustness. Drawing from episodic memory, the ability to retrieve specific experiences with contexts(Tulving et al., 1972), Zhang et al. (2025a) enable LLMs to learn from episodic experiences for improved planning, Zhang et al. (2025c) integrates episodic memory with semantic memory for LLM personalization, while Chaudhury et al. (2025) introduce episodic attention for processing long contexts.
Reinforcement Learning from Human Feedback (RLHF) A classic behavioral theory, the Operant Conditioning theory posits that behaviors are systematically strengthened or weakened by the consequences (rewards or punishments) that immediately follow them (Skinner, 1957; Thorndike, 1898) principles of reinforcement learning align closely with this psychological framework, particularly in the post-training phase of LLM development, where RLHF explicitly operationalizes Operant Conditioning theory to align model behaviors with human values and preferences. Through repeated feedback, the model gradually adapts to favor outputs that yield higher reward signals—a process akin, in a loose analogy, to Thorndike’s Law of Effect, which describes how behaviors followed by satisfying outcomes tend to recur. While the underlying mechanism is driven by reward optimization algorithms rather than psychological intent, the conceptual resemblance highlights how reinforcement strategies can shape model outputs (Lambert et al., 2023). During RLHF, the model generates responses, and a learned reward function $R(x)$ assigns scores to outputs $x$, guiding subsequent policy updates. For instance, Ouyang et al. (2022) train InstructGPT using Proximal Policy Optimization (Schulman et al., 2017), rewarding responses preferred by humans and penalizing less desirable ones. Foundational frameworks (Christiano et al., 2017; Stiennon et al., 2022; Sutton and Barto, 2018)explicitly translating human judgments into reward signals, operationalizing the insights of Operant Conditioning. More recent work incorporates human cognitive biases (Siththaranjan et al., 2024; Wadi and Fredette, 2025) and personalizes reward functions for individual values (Poddar et al., 2024) and awareness (Liu et al., 2025a). These developments illustrate how Operant Conditioning remains central to aligning LLMs with nuanced human values. While our survey focuses on psychological dimensions, a technical overview of RLHF methods is provided in Appendix LABEL:appendix:rlhf.
5 Evaluation and Application
Psychology offers tools for both assessing and enhancing model behavior in the evaluation and application stage. We review three key areas of challenges where psychology can inform LLM development: (1) evaluating emergent capabilities such as reasoning, (2) improving task performance in domains involving human cognition, and (3) designing socially aware, multi-agent systems.
5.1 Benchmarks and Capability Assessment
Evaluating LLMs with psychologically grounded metrics offers a deeper window into their real-world viability. By mapping classic theories onto benchmarks that probe model responses under diverse, human-like scenarios, researchers can move beyond surface-level performance measures, revealing emergent model behavior and illuminating strengths, blind spots, and opportunities to refine LLM training and alignment practices.
5.1.1 Social Reasoning and Intelligence
Social intelligence is vital for LLMs that navigate human contexts, enabling the interpretation of implicit cues, adaptation to social norms, and authentic interaction – defining advanced AI beyond mere text prediction. As LLMs increasingly mediate communication, their grasp of social dynamics becomes pivotal for both efficacy and safety.
Notably, Theory of Mind (ToM) offers a framework for evaluating how individuals understand and attribute mental states – such as beliefs, desires, and intentions – to others. By measuring LLMs’ capacity to reason beliefs, researchers can assess core social intelligence. Recent benchmarks probe different facets of ToM: Chen et al. (2024c); Xu et al. (2024a) provide relatively comprehensive coverage, while others test ToM on higher-order reasoning (Wu et al., 2023), interactions (Kim et al., 2023), white lies (Liu et al., 2025d), cognitive appraisal (Yeo and Jaidka, 2025), and temporally evolving mental states (Xiao et al., 2025). Extending the efforts to spoken dialogues, Soubki et al. (2024) reveal lingering gaps between LLM and human performance. Surveys (Chen et al., 2025b; Ma et al., 2023; Sap et al., 2022; Soubki and Rambow, 2025) consolidate methods and underscore the challenges of robust ToM-based evaluations.
Beyond individual cognition, social influence theories like Conformity Theories (Asch, 2016), capture how group pressure shapes individual judgments. Recent work tests LLM-based agents’ collaboration and bias dynamics under these principles (Choi et al., 2025; Jin et al., 2024b; Wu et al., 2025d; Zhang et al., 2024c; Zhu et al., 2025), bridging individual and group-level cognition.
Emotion is another pillar of social intelligence. Ekman’s Basic Emotion Theory (Ekman, 1992) identifies six universal emotions, often used as labels, while Dimensional Models like the Circumplex Model conceptualize emotions along valence and arousal (Gong et al., 2024; Morrill et al., 2024). LLMs advance on emotion recognition, benefiting dialogue and sentiment tasks (Sabour et al., 2024; Wu et al., 2024c, d; Zhang et al., 2024e).
These efforts collectively demonstrate both progress and limitations in LLMs’ social cognition, establishing benchmarks against which future developments can be measured.
5.1.2 Language Proficiency
Recent work adopts psycholinguistic assessments, originally designed for humans, to test LLMs’ language proficiency. These experiments probe a wide range of linguistic domains: morphology (Anh et al., 2024), syntax (Amouyal et al., 2025; An et al., 2025; Hale and Stanojević, 2024; Kandra et al., 2025; Li and Hao, 2025; Liu et al., 2024b), phonology (Goriely and Buttery, 2025; Jang et al., 2025), semantics (Bavaresco and Fernández, 2025; Duan et al., 2025; Hayashi, 2025) and their combinations (Ma et al., 2025a; Miaschi et al., 2024; Zhou et al., 2025a).
Although LLMs exhibit comparable performance to human speakers on many psycholinguistic tasks, the underlying processing mechanism they rely on may seem different from humans (Lee et al., 2024; Pedrotti et al., 2025). Human language acquisition is often characterized by the Poverty of the Stimulus, children acquire complex grammar from relatively little input (Chomsky, 1980), whereas LLMs typically require developmentally implausible amounts of linguistic data to learn morphological rules. On the other hand, some evidence suggests that the learning patterns of LLMs mirror aspects of human language acquisition (Liu et al., 2024b; Zhou et al., 2025c).
Several studies have explored the pragmatic abilities of LLMs, motivated by the close link between language and broader cognitive functions in humans. Grice (1975)’s Theory of conversational implicature posits that utterance interpretation depends on both literal content and surrounding context. Researchers (Bender and Koller, 2020; Gubelmann, 2024) have contrasting perspectives on LLMs with respect to the Harnad (1990)’s Symbol Grounding Problem, i.e. linguistic symbols must be grounded in sensorimotor interactions to be meaningful. Failures of LLMs in pragmatic and semantic tasks (He et al., 2025; Junker et al., 2025; Kibria et al., 2024; Zeng et al., 2025; Ma et al., 2025a), as well as their neuron patterns (He et al., 2025; Wu et al., 2024b), point to limitations beyond pure linguistic knowledge, which potentially parallel human higher-level cognitive processes.
5.1.3 Memory and Cognitive Evaluation
Assessing memory and cognition is crucial given LLMs’ limited capacity and risk of catastrophic forgetting. Memory is measured on parametric knowledge (Li et al., 2023), n-back tasks (Zhang et al., 2024a), capacity (Timkey and Linzen, 2023) and cognitive load (Fu et al., 2025; Xu et al., 2024b; Zeng et al., 2024b). Meanwhile, cognitive development is assessed through cognitive maturity (Laverghetta Jr. and Licato, 2022; Wang et al., 2025b), word acquisition (Chang and Bergen, 2022), subjective similarity (Malloy et al., 2024), reasoning strategies (Mondorf and Plank, 2024; Yuan et al., 2023; Ying et al., 2024), zone of proximal development (Cui and Sachan, 2025) and perception (Jung et al., 2024; Zhou et al., 2025b).
5.1.4 Personality Capability
Personality consistency examines how stably LLMs maintain traits across contexts. Frisch and Giulianelli (2024) show LLMs with asymmetric profiles vary in Big Five traits, while Amidei et al. (2025) find language switching alters GPT-4o’s Eysenck Personality Questionnaire Revised traits, underscoring challenges in perserving consistency. Parallel research examines how LLMs display and control personality traits. Jiang et al. (2024) show LLMs express distinct traits labeled by human evaluators. Mao et al. (2024) reveals difficulties in alignment for Neuroticism, Extraversion and Agreeableness. Han et al. (2025); Lee et al. (2025b); Li et al. (2025c, d) assess and improve consistency through alignment with psychometrical training data, while Hu and Collier (2024) find persona-based prompting improves annotation accuracy.
5.1.5 Bias and Ethics Evaluation
Evaluating biases and ethical risks is crucial for responsible AI that avoids reinforcing harmful social patterns. As LLMs increasingly shape public discourse, thorough assessments are essential to prevent discriminatory outputs and promote equitable benefits across diverse communities. Recent work tests LLMs on gender (Oba et al., 2024; Zhao et al., 2024), broader social biases (Abolghasemi et al., 2025; Huang et al., 2025a; Lee et al., 2023; Muti et al., 2025; Nozza et al., 2022; Shin et al., 2024; Zhao et al., 2025), toxic content (Gehman et al., 2020; Huang et al., 2025c; Hui et al., 2024a; Luong et al., 2024), and harmful stereotypes (Grigoreva et al., 2024; Huang and Xiong, 2024; Hui et al., 2024b; Shrawgi et al., 2024), establishing benchmarks across cultures and languages. Evidence also suggests that LLMs replicate social identity biases, mirroring human tendencies toward ingroup favoritism and outgroup hostility (Borah et al., 2025; Dong et al., 2024a; Hu et al., 2025b; Wang et al., 2025c) – patterns central to social identity theory, which posits that group membership shapes self-concept and intergroup behavior (Tajfel, 1979).
5.2 Task Performance Enhancement
Building on the benchmarks, we review how psychological insights are used to improve LLMs performance on complex reasoning and enrich dialogue, which illustrate how psychology improves capabilities and alignment across applications.
5.2.1 Reasoning Enhancement
LLMs often struggle with complex reasoning: social inference (Liu et al., 2024a), logical errors (McKenna et al., 2023; Turpin et al., 2023), hallucinations (Ai et al., 2024a; Huang et al., 2025b), and multi-step planning (Wang et al., 2024a). Researchers address these issues by implementing analogous cognitive mechanisms. For instance, Dual-process theory, a social cognition framework, distinguishes between fast (System 1) and slow (System 2) reasoning (Kahneman, 2011), offers a blueprint for LLM improvement. Chain-of-thought prompting (Wei et al., 2022) operationalizes System 2 via intermediate steps, while DynaThink (Pan et al., 2024) dynamically selects rapid or thorough inference. Tree of Thoughts (Yao et al., 2024) further explores multiple reasoning paths concurrently. Yang et al. (2025) combine separate verifier as System 2. More recent applications includes hallucination mitigation (Cheng et al., 2025c), real-time human-AI collaboration (Zhang et al., 2025b), multi-hop QA (Cheng et al., 2025a), emotion consistency (Wei et al., 2025), decoder-level LLMs merging (Hu et al., 2025a) and cost-efficient RL framework (Cheng et al., 2025b).
Similarly, Self-reflection and Meta-cognition, introspection focused on the self-concept (Flavell, 1979; Phillips, 2020), has guided LLM enhancements in hallucination mitigation (Ji et al., 2023; Lu et al., 2025), translation (Chen et al., 2024a; Wang et al., 2024e), tool use (Li et al., 2025b), question-answering (Kassner et al., 2023; Li et al., 2024a; Zhang et al., 2024f), retrieval-augmented-generation(RAG) (Asai et al., 2024; Zhou et al., 2024) and math reasoning (Zhang et al., 2024f). Approaches include iterative self-assessment (Ji et al., 2023; Wu et al., 2025c; Yan et al., 2024), task decomposition (Wang et al., 2024e; Zhang et al., 2024d), self-training (Dou et al., 2024), in context learning (Ji-An et al., 2025), and confidence-tuned reward functions (Xu et al., 2024c). Moreover, ToM adaptations boost LLMs’ interpersonal reasoning, aiding missing knowledge (Bortoletto et al., 2024), common ground alignment (Qiu et al., 2024), and cognitive modeling (Wu et al., 2024a).
Beyond social reasoning, perception, attention, and memory support coherence and retrieval. Kojima et al. (2022) uses “think step by step” prompts for top-down reasoning. Li et al. (2025a) employed Piaget’s constructivist theory on a systematic memory system for reading comprehension. Chen et al. (2025a); Maharaj et al. (2023); Yu et al. (2022) leverages selective attention and working memory to detect hallucinations and extract relation. Zhu et al. (2024) employs recitation for retrieval, and Park and Bak (2024) introduce short/long-term memory modules. Chi et al. (2023); Diao et al. (2025); Wang et al. (2024c) improve reasoning via symbolic, adaptive and working memory structures. Lastly, hippocampal indexing theory (Teyler and DiScenna, 1986), viewing the hippocampus as a pointer to neocortical memory, informs multi-step reasoning with external knowledge (Fang et al., 2025; Gutierrez et al., 2024) and counterfactual reasoning (Miao et al., 2024a).
5.2.2 Dialogue Understanding and Generation
In dialogue understanding, personality psychology aids trait-based inferences from user interactions. NLP research has explored dynamic ways to measure personality beyond structured tests. The Myers–Briggs Type Indicator (MBTI), a self-report questionnaire that makes pseudo-scientific claims to categorize individuals into 16 distinct personality types, remains popular (Rao et al., 2023; Yang et al., 2023), while PsychoGAT (Yang et al., 2024) gamifies MBTI, and PADO (Yeo et al., 2025) adopts a Big Five-based multi-agent approach. Beyond assessments, traits guide dialogue generation: Huang and Hadfi (2024) show higher agreeability improves negotiation, while Cheng et al. (2023) reveal social and racial biases in persona creation, raising representational concerns.
Dialogue generation research further incorporates personality to improve coherence, empathy, and consistency. Dan et al. (2025) enforced trait-consistent generation via Big Five trait specialization loss. Joshi et al. (2025) augmented personas with rationales generated by Big Five-guided scaffolds. Chen et al. (2025c); Pal et al. (2025) leveraged Reddit-based journal entries to model Big Five traits in large-scale dialogue datasets. Big Five-aligned agents also improve on text based games (Lim et al., 2025), decision making under risk (Hartley et al., 2025) and code generation (Guo et al., 2025). Other efforts improve persona consistency without referencing explicit psychological theory (Takayama et al., 2025; Wu et al., 2025b). Similarly, personality is used to improve truthfulness, consistency, and context-aware generation, as further detailed in Appendix LABEL:appendix:persona_dialogue. These approaches support personality alignment but lack grounding in deeper psychological theory.
5.3 Collaborative, Multi-Agent Frameworks
Beyond task-specific capabilities, the surge in multi-agent LLM frameworks reflects a growing emphasis on collaborative decision-making, where modeling social dynamics is crucial. Social and personality psychology theories offer insights to design agent interaction, negotiation, and consensus, guiding more socially intelligent LLM systems.
Social Influence
Persuasion models (Petty and Cacioppo, 2012) illustrate how central/peripheral routes shape attitudes in collaborative settings. Leveraging this, Gollapalli and Ng (2025) merges persuasive dialog acts with RL, Modzelewski et al. (2025) infuses persuasion knowledge into CoT, Furumai et al. (2024) combines LLM strategies and retrieval, Jin et al. (2024a); Qin et al. (2024) emphasize credibility-aware generation, and Hwang et al. (2025); Tan et al. (2025); Zeng et al. (2024a) uncovers LLMs’ vulnerabilities. Multi-agent research simulates personality-driven (Huang and Hadfi, 2024; Hu et al., 2025c; Liu et al., 2025b) and cultural-driven negotiation (Aghazadeh and Kovashka, 2025; Ma et al., 2025b), boosts truthfulness via structured debates (Khan et al., 2024), and curates argument-strength datasets (Rescala et al., 2024).
Social Cognition
ToM complements social influence by enabling agents to grasp others’ mental states. Some integrate perspective-taking (Lică et al., 2025), belief tracking (Sclar et al., 2023) and coordination (Wang et al., 2022; Sclar et al., 2022), while others refine ToM via task decomposition and recursive simulation (Jung et al., 2024; Sarangi et al., 2025; Wilf et al., 2024).
Role-Play and Multi-Agent Simulation
Recent work on persona-driven LLM agents focuses on simulating diverse perspectives, persona alignment, and socially intelligent interactions. Han et al. (2024) introduces Big Five-based extraversion, Castricato et al. (2025) presents 1,586 synthetic personas, and Wu et al. (2025a) releases a benchmark with 40K multi-turn dialogues. Agents also model opinion dynamics (Wang et al., 2025a) and evaluate social intelligence (Chen et al., 2024b), with RoleLLM (Wang et al., 2024b), Character100 (Wang et al., 2024d), and persona-aware graph transformers (Mahajan and Shaikh, 2024) further supporting multi-party simulations. Lastly, Kumarage et al. (2025) and Chen et al. (2026) simulate social engineering attacks with LLM and Audio LM agents of varied traits, highlighting how psychological traits shape user vulnerability and the need for more robust defenses against personalized manipulation.
6 Trends and Discussion
6.1 How Does Current LLM Development Harness Psychological Theories?
From the review from Section 1-5, we observe psychological theories have been incorporated into LLM development in stage-specific ways, with uneven coverage across theoretical domains. Figure 1 maps this integration across stages.
In early stages, preprocessing and pretraining, developmental psychology is often referenced. Its emphasis on staged knowledge acquisition aligns with curriculum learning and progressive data exposure, mirroring human developmental trajectories. In post-training, especially RLHF, behavioral psychology ideas are most prominent. Conditioning, reinforcement schedules, and reward design are commonly used to guide model alignment with human preferences. In evaluation and application, theories from social psychology, personality psychology and psycholinguistics are commonly cited, reflecting a focus on interaction patterns, user modeling, and linguistic variation – areas traditionally explored within these sub-fields. Their prominence in later stages aligns with their emphasis on human-centered communication. Cognitive psychology appears across all stages, particularly in modeling internal mechanisms such as reasoning, memory, and attention. Its breadth makes it a foundational influence, though less visible in agentic interaction settings.
The observed unevenness in integration reflects, perhaps a gap, but more probably a functional alignment – some domains are naturally better suited for certain stages of LLM development. Meanwhile, these trends expose under-explored opportunities, motivating the RQs that follow.
6.2 What Untapped Psychological Insights Could Advance LLM Development?
Although psychological theory is increasingly applied in LLM research, its use remains simplified and uneven. As shown in Tables LABEL:tab:psych_theories_dev_cog_beh, LABEL:tab:psych_theories_soc, and LABEL:tab:psych_theories_per_ling, many theories are under-utilized despite their potential to improve model behavior and interpretability. Below, we outline theories in four key areas that deserve greater attention in future LLM research.
Social psychology remains underutilized in areas like group dynamics and self and identity, limiting personalization, adaptability, and inclusivity. Prompting LLMs to adopt specific social identities can reduce bias (Dong et al., 2024a) and mirror human-like ingroup favoritism (Hu et al., 2025b). Incorporating social identity frameworks could enhance user alignment in identity-sensitive contexts (Chen et al., 2020). Likewise, while bias detection is common, classic social influence theories (e.g., conformity, obedience) and attitude change theories (e.g., balance theory, cognitive dissonance) are rarely applied to interaction dynamics or bias mitigation, despite their relevance to ethical and socially adaptive behavior. Additionally, malicious actors leveraging social influence can severely undermine trust in digital spaces (Ai et al., 2024b; Liu et al., 2025c; Zeng et al., 2024a), highlighting the potential of constructs like inoculation theory to proactively guard against manipulative strategies.
Behavioral psychology inspires RLHF, yet key concepts like partial reinforcement, which improves behavior persistence (Ferster, 1966; Jensen, 1961), and shaping, which supports gradual learning through successive approximations (Love et al., 2009), are overlooked. Current RLHF relies on uniform rewards, yet behavioral theory warns that flawed rewards can lead to reward hacking. Adding reward variability may reduce premature convergence and improve alignment with human intent (Amodei et al., 2016; Dayan and Daw, 2008).
Personality Psychology use focuses on Trait Theory, overlooking developmental theories that explain how individual traits emerge, evolve, and adapt across contexts. These developmental models could enable more coherent and interpretable personality representations, offering a deeper alternative to static prompt-based personas.
Cognitive psychology remains underused, particularly Schema Theory, which holds that humans store knowledge as schemas formed through repeated experience (Anderson and Pearson, 1984), guiding inference, memory, and learning. Recent work explores schema-inspired methods for compressing user histories and modeling knowledge activation cycles (Panagoulias et al., 2024; Xia et al., 2024). Further integration may improve long-term context handling and generalization.
6.3 What Debates Loom at the NLP– Psychology Intersection, and Where Next?
A recurring question is whether human psychology can be directly mapped to LLMs without distortion (Löhn et al., 2024). Below, we highlight key controversies at this boundary; see Appendix LABEL:appendix:debates for an extended discussion. These challenges motivate new recommendations and highlight open directions for cross-disciplinary exploration.
Terminology Mismatches
A core tension is the mismatch between psychological terminology and their NLP usage. For example, attention in psychology refers to selective mental focus, but in transformers it is a token weighting mechanism without cognitive awareness (Lindsay, 2020), leading to misleading attributions of intentionality. Similarly, memory in psychology entails structured encoding and recall, whereas in LLMs it typically refers to context windows or parameters.
Such anthropomorphic language is increasingly prevalent and shapes public and scholarly assumptions about LLMs, as recent studies show rising human-like descriptors (Ibrahim and Cheng, 2025). This calls for disentangling metaphor from mechanism through a precise cross-disciplinary lexicon, preventing both oversimplification and over-anthropomorphization – an underexplored but crucial research challenge.
Theoretical Discrepancies
Beyond terminology, deeper theoretical mismatches arise when the NLP community adopts outdated or disputed concepts from psychology. For instance, predictive coding (Rao and Ballard, 1999) is used to analogize LLMs’ next-token prediction, although current research emphasizes hierarchical, multi-scale brain mechanisms (Antonello and Huth, 2024; Caucheteux et al., 2023). Likewise, folk-psychological typologies like MBTI persist in LLM applications despite its criticized validity and reliability (McCrae and Costa Jr, 1989; Pittenger, 1993). Wagner et al. (2025) positions that ToM involves first deciding depth of mentalizing and then applying reasoning accordingly, yet most works focus only on the latter. Working memory (Baddeley and Hitch, 1974) illustrates another gap: LLM "memory" modules (Li et al., 2023; Kang et al., 2024) do not replicate human constraints, prompting questions about whether AI should emulate human cognitive limits or exceed them for performance gains. Behavioral psychology faces similar critiques (Flavell et al., 2022; Miller, 2003), as RLHF often focuses on reward optimization (Ouyang et al., 2022; Rafailov et al., 2023; Ramesh et al., 2024), neglecting internal states and risking reward hacking (Krakovna, 2020; Skalse et al., 2022). Broader debates remain over whether LLMs truly “understand” language or function as “stochastic parrots” (Ambridge and Blything, 2024; Park et al., 2024).
In response, we recommend refining how psychological theories are mapped into computational models, replacing outdated constructs with supported frameworks, exploring whether human-like constraints aid interpretability, and designing evaluations that track both outputs and internal states. Sustained collaboration between computational and psychological sciences is essential for robust and theory-aligned LLMs.
Evaluation and Validity Debates
Another major debate is how we evaluate LLM “psychological” abilities – whether current tests really measure what they claim. For instance, GPT-4 solves around 75% of false-belief tasks, matching a 6-year-old’s performance (Kosinski, 2024; Strachan et al., 2024); some see emergent ToM-like reasoning (Kosinski, 2024), but others argue it may be pattern matching (Strachan et al., 2024), noting that minor prompt changes can derail results (Shapira et al., 2024). This calls for more theory-grounded evaluation and clearer definitions.
A parallel controversy involves personality. Some studies find stable simulated personality traits (Sorokovikova et al., 2024; Huang et al., 2024), while others reveal variability under different prompt conditions (Gupta et al., 2024; Shu et al., 2024), raising questions about inherent vs. mimicked personas Tseng et al. (2024). These debates underscore the need for a systematic, theory-driven framework beyond surface metrics, guiding more faithful replication of human cognition and behavior in LLMs.
These debates also echo concerns raised within Psychologists’ Perspectives on LLMs, where scholars are critically examining the promises and limitations of applying human psychological constructs to LLMs. Psychologists are engaging LLMs both as tools for research and as objects of study, but they caution against straightforwardly mapping human psychological constructs onto models trained via next-token prediction on disembodied text. Method papers flag opportunities and pitfalls for using LLMs as participants, judges, or simulators, urging stronger validity checks and domain knowledge when importing instruments or constructs Abdurahman et al. (2024b). At the theoretical level, scholars argue we should be precise about what LLMs are meant to model and avoid inferring humanlike learning or cognition from surface similarities Blank (2023); Cuskley et al. (2024). Empirically, comparative studies probe where model behavior diverges from humans on socially loaded tasks: LLMs can pass certain ToM-style probes while still differing on harder variants Strachan et al. (2024); Jones et al. (2024), and they sometimes amplify cognitive biases such as omission bias in moral judgment Cheung et al. (2025). Together, these works show an active psychology literature investigating LLMs’ promise and limits, and underscores why direct transfer of human psychological theory to LLMs requires careful construct validity, task design, and interpretation.
7 Conclusions
We systematically review how psychology can ground LLM innovation in both past and future across six key subfields. We examine how psychological theories inform each stage of LLM development, revealing both meaningful connections across domains and critical points of tension, which we explore through discussion to help bridge interdisciplinary gaps. We hope this survey sparks reflection, and inspires future work to continue integrating psychological perspectives into NLP in meaningful and impactful ways.
Limitations
Our review primarily focuses on literature within NLP, particularly in how personality is modeled, evaluated, and leveraged in LLMs. As a result, we do not extensively cover research from psychology and cognitive sciences that might offer deeper theoretical insights into human-like behaviors in AI. This limitation may exclude valuable methodologies or perspectives that could enhance personality evaluation frameworks for LLMs. We encourage future surveys to integrate findings from psychology and linguistics to bridge theoretical foundations with computational approaches, fostering a more comprehensive understanding of personality in AI systems. Some perspectives (e.g., embodied cognition, constructivism) cut across our categories and appear sporadically in current LLM work; we note these as cross-cutting threads and encourage future surveys to map them more fully across stages."
While our survey advocates for a deeper integration of psychology into LLM design, we also caution against the ethical risks posed by overuse or misapplication of psychological principles. A concrete example is operant conditioning (Skinner, 1957), which describes how behavior can be shaped by consequences. Applied to LLMs, these mechanisms can be beneficial in contexts like language learning or motivation. For instance, through timely, gratifying feedback to reinforce engagement. However, reinforcement schedules such as variable ratio or interval rewards may unintentionally condition users to engage compulsively, raising the risk of manipulative design. This presents a key ethical limitation: distinguishing between genuinely supportive interactions and those that encourage excessive use is inherently difficult. To address this, we emphasize the need for transparent disclosure of reinforcement mechanisms and the establishment of clear ethical guidelines by professional communities. These safeguards are essential to ensure that psychological insights enhance user well-being without enabling exploitative practices.
We also recognize longstanding critiques of psychology, for example, its emphasis on individual-level constructs, WEIRD sampling (overreliance on participants from Western, Educated, Industrialized, Rich, and Democratic societies in behavioral and social science research), and under-representation of social/structural forces. Our survey foregrounds psychology because many current LLM papers explicitly reference it; however, insights from linguistics and sociology are equally important for understanding language as a fundamentally social phenomenon. Future surveys that braid these traditions more tightly would provide a more complete view.
An important adjacent question is what LLMs do psychologically to users (e.g., persuasion, behavior change, support). We flag this as a crucial direction for future surveys that integrate psychology with HCI and communication research to evaluate user-level outcomes alongside model-level capabilities.
Acknowledgments
This work is in part supported by the funds provided by the National Science Foundation and by DoD OUSD (R&E) under Cooperative Agreement PHY-2229929 (The NSF AI Institute for Artificial and Natural Intelligence). The views, opinions and/or findings expressed are those of the authors and should not be interpreted as representing the official views or policies of the National Science Foundation or the U.S. Government.
Appendix A Psychology Theories
\__nicematrix_patch_booktabs:\__nicematrix_revert_colortbl: